A train travels on a straight track passing a signal A at 20m/s. It accelerates uniformly at 2m/s and reaches signal B 100 m further than A. At B the velocity of the train in m/s is?
1
Expert's answer
2020-11-17T07:12:45-0500

Initial velocity "u=20m\/s"


Acceleration "a= 2m\/s^2"


Distance "S= 100m"


Final velocity ( Velocity at pole B) ("v" ) :


"v^2 = u^2+2aS"


"v^2=20^2+2\times2\times100=800"


"v=28.28m\/s"